my equation has a trianglewith the hypotenuse at 10 centimeters the adjacent side is 3x centimeters and my opposite is x centimeters. I know that the pythagoreum theroreum is A squared + B squared = C squared and I got to the point where X centimeters squared + 3X centimeters squared = 100 centimeters could you help me please?
Answer by checkley71(8403)

x^2+3x^2=100
4x^2=100
x^2=100/4
x^2=25
x=sqrt25
x=5 answer,
proof:
25+3*25=100
25+75=100
100=100